# Public-private funding models in open source software development: A case study on scikit-learn

### v3

@@ Abstract
Governments are increasingly funding open source software (OSS) development to address concerns regarding software security, digital sovereignty, and national competitiveness in science and innovation. While announcements of governmental funding are generally well-received by OSS developers, we still have a limited understanding of how they evaluate the relative benefits and drawbacks of such funding compared to other types of funding. This paper explores this question through a case study on scikit-learn, a Python library for machine learning, whose funding combines research grants, commercial sponsorship, community donations, and a €32 million grant from France’s artificial intelligence strategy. Through 25 interviews with scikit-learn’s maintainers and funders, this study makes two key contributions to research and practice. First, the study contributes novel findings about the design and implementation of a public-private funding model in an OSS project. It sheds light on the respective roles that public and private funders have played in supporting scikit-learn, and the processes and governance mechanisms employed by the maintainers to balance their funders’ diverse interests and to safeguard community interests. Second, it offers practical recommendations. For OSS developer communities, it illustrates the benefits of a diversified funding model for balancing the merits and drawbacks of different funding sources and mitigating dependence on single funders. For companies, it serves as a reminder that sponsoring developers or OSS projects can significantly help maintainers, who often struggle with limited resources and towering workloads. For governments, it emphasises the importance of funding the maintenance of existing OSS in addition to funding the development of new software or features. The paper concludes with suggestions for future research.
Keywords Open source software  $\cdot$ OSS funding models  $\cdot$ machine learning  $\cdot$ scikit-learn  $\cdot$ OSS sustainability
@@ 1 Introduction
Open source software (OSS) has moved up the priority list of governments in light of concerns about digital sovereignty [1], software security [2], and national competitiveness in science and innovation [3], amongst others. For example, the discovery of the Log4Shell vulnerability in November 2021, described as the “worst security problem in a generation” [4], prompted swift action from governments. In January 2022, the White House convened cross-sector stakeholders, including government agencies, Big Technology companies, and OSS foundations, to discuss the collective responsibility to improve the security of OSS, recognising its indispensable role as digital infrastructure whose maintenance largely depends on volunteers [5]. In June 2022, the newly formed European Working Team on the Digital Commons, comprising representatives from 19 member states of the European Union (EU) and the European Commission (EC), published its inaugural report, Towards a Sovereign Digital Infrastructure of Commons, highlighting the importance of funding OSS for bolstering the digital sovereignty of Europe [6]. Some months later, in October 2022, the German government established the Sovereign Tech Fund (STF), allocating a budget of €11.5 million for its first year alone to fund the development, improvement, and maintenance of open digital infrastructure to support digital sovereignty, software security, innovation, and digital democracy [7]. These developments underscore the growing recognition of OSS as digital public goods and the public sector’s responsibility to participate in funding its development [8].
@@
The increasing interest and involvement of governments in funding OSS development has been generally well-received by OSS developers. However, the question of how OSS developers view the relative benefits and drawbacks of governmental funding compared to other sources of funding remains unexplored. This paper addresses this timely issue through a case study on scikit-learn, a popular Python library for machine learning (ML), which is described as “the Swiss army knife of ML” due to its widespread use in research and industry [9]. scikit-learn is developed by a team of core maintainers, who are primarily based at Inria, the French Institute for Research in Computer Science and Automation, and a global community of volunteers. Its sustainability as a community-led OSS project in the industry-dominated field of artificial intelligence (AI) [10] is in part attributed to its mixed funding model, including research grants, commercial sponsorship, and community donations, as well as the institutional backing of Inria since 2010, which have provided the means to employ full-time maintainers, to provide them with office space and IT resources, and to organise community events, amongst others. November 2021 was a turning point for scikit-learn. Recognising the significance of scikit-learn and OSS in general in AI research and development (R&D), the French government announced a €32 million grant in its AI strategy to support the expansion of scikit-learn into a comprehensive suite of open source tools for ML and AI for the benefit of researchers and developers in France and beyond [11].
Through 25 semi-structured interviews with the scikit-learn maintainers and their public and private funders, this paper explores the historical role of mixed public-private funding in supporting the scikit-learn project. The paper makes two key contributions with implications for research and practice. First, the study contributes novel empirical findings about the design and implementation of a public-private funding model in an OSS project. The analysis sheds light on the respective roles that public and private funders have played in supporting scikit-learn and how its maintainers have balanced and when necessary negotiated their funders’ diverse interests. Second, the study provides practical insights for practitioners. For OSS developer communities, it shows that a diversified funding model can act as an effective means for balancing the merits and drawbacks of different funding sources, as well as for containing the influence of any single funder, whether they are a company, a government body, or other. For companies, it is a timely reminder that by sponsoring developers or an OSS project, they can make a substantial difference to OSS maintainers, who struggle with towering workloads and limited resources. For governments, it underscores the importance of funding maintenance of existing OSS and refraining from the novelty bias of funding new software or features.
@@
The paper has the following structure. First, it discusses prior work on the role of funding, amongst other mechanisms, in supporting the sustainability of OSS projects. Second, it presents the study design and the case of scikit-learn. Then, it presents the results and discusses their implications for research and practice. The paper concludes with a discussion of the key recommendations to practitioners and suggestions for future research on OSS funding models.
@@ 2.1 The role of funding in OSS sustainability
The sustainability of OSS projects is a multi-faceted problem. With the increasing recognition of OSS as digital infrastructure [12, 13, 14], an extensive literature has examined the reasons for OSS project failures and successes. A range of factors can contribute to the sustainability of OSS projects [15], from the steady inflow of new contributors [16, 17, 18, 19, 20] to a number of maintainer “best practices,” including timely responses [21], mentoring [22, 20], and task recommendation [23]. Others point to the role that foundations play as “boundary organisations” [24], which enable collaboration between diverse contributors, including volunteers and for-profit companies, who collectively test, develop, and maintain OSS [25, 26, 27]. Conversely, prior work shows that a host of social, economic, and technological factors can cause OSS projects to fail (or “deprecate”), including usurpation by a competitor, a lack of time or interest amongst maintainers, conflicts amongst maintainers, or the obsolescence of the software, amongst others [28].
@@
The financial health of an OSS project is an important variable for its sustainability, concerning if or to what extent its maintenance is funded or reliant on volunteers [29]. However, money has long been a contentious subject amongst OSS developers. On the one hand, many developers take pride in the origins of the OSS movement as a social movement [30], described as a “programmers’ paradise”, comprising “geeks”, “hackers”, and “hobbyists,” who do not work for money [31]. Many developers cite their political ideals [32], altruism [33], or their passion for working on hard problems [34] as intrinsic incentives for contributing to OSS. Many complain that “money ruins everything” [35]. On the other hand, the OSS ecosystem is awash with unfunded projects and overstretched volunteers, who struggle to keep up with a spectrum of “visible” and “invisible” tasks required to keep projects running [29, 36, 21].
@@
The public disclosure of vulnerabilities in widely used OSS libraries, such as Heartbleed in April 2014 or Log4Shell in December 2021, which was described as the “worst security problem in a generation” [37], captured the attention of industry and governments alike. The latter brought to light the systemic issues of free-riding and under-investment in OSS as digital public infrastructure, and the need for cross-sector stakeholders, including the public and private sectors, to financially support the developer communities who build and maintain this infrastructure [8]. Indeed it has been estimated that OSS is present in 96% of codebases [38] and constitutes up to 90% of software stacks [39]. As Benjamin Bikinbine has argued, there is an urgent need “not just [for] investment in institutions, organisations, technologies, or innovations, but long-term and sustainable investment in the true source of their value: people” [40].
2.2 Funding models in OSS development
Funding models for OSS have been evolving in recent years, with individuals, public bodies, companies, and philanthropies playing a part in financially supporting OSS development and maintenance. Individuals—who may be OSS developers themselves or simply users of OSS—fund OSS developers by making micro-donations. For instance, the GitHub Sponsors scheme, launched in 2019, allows users to make one-off or recurring payments to individuals or organisations with sponsored profiles [41]. According to prior work, common incentives for donations via GitHub Sponsors include expressing appreciation, recognising a developer’s work, and encouraging future contributions [42]. By comparison, common incentives for maintainers to accept donations are to receive recognition or to provide a side income for their OSS work; while those that do not participate state that they do not need to be sponsored or they do not contribute to open source for money [42]. Other work shows that the amount that individuals donate is associated with the length of their use of an OSS [43]. The impact of micro-donations is still contested. On the one hand, they have been shown to shorten response times to issues [44, 45] and to increase maintenance-related activities [46]. On the other hand, the impact appears trivial [47] and donations tend to be ad-hoc, small, or legally complicated to receive [48].
@@ 2.2 Funding models in OSS development
The private sector has been the largest funder of OSS to date, funding OSS developers and projects in both direct and indirect ways. The impact of commercial investments in OSS are gauged to be significant; according to one estimate, companies in the EU invested around €1 billion in OSS development in 2018, which generated €65-95 billion for EU GDP that year [49]. Direct funding methods include sponsorship of OSS consortia or foundations, which operate on a fee-paying membership model [50]. Recently, companies have begun to establish FOSS Contributor Funds, which allow employees to nominate and vote on OSS projects that should receive funding [51]. By involving individual developers in the funding decision-making process, these funds facilitate a more democratic way for companies to support the OSS that they use in their products or services [52]. A shortcoming of these novel funds is that funding awards are typically limited to \$10,000 over a one-year time frame, which does not amount to adequate support for OSS projects in the long run. Alyssa Wright from Bloomberg’s FOSS Contributor Fund ([53]) has acknowledged this shortcoming, arguing that “true sustainability comes from a multi-sided commitment” and “while financial support is important, real engagement through contribution and dialogue is as crucial to the ecosystem’s survival as writing a check.”
@@
Other ways companies support OSS projects is by sponsoring OSS developers or by allowing their employees to contribute to OSS at work [54, 55]. Sponsoring developers is a recognised strategy that companies employ to influence projects [56, 57] and to improve their reputation as OSS patrons, which in turn is considered to aid the recruitment of OSS developers [58, 59, 60]. In the mid-2000s, it was estimated that 40% of OSS contributors were paid [61]. We should expect this number to be much higher today, with companies like Google, Microsoft, and Amazon accounting for a rapidly growing number of OSS developers and contributions to OSS projects [62]. In addition, companies invest in projects through donations and joining project steering committees [63]. While OSS projects benefit from commercial support, maintainers seek to limit the influence or even the dominance of companies in projects [64, 65]. To manage the influence of companies, OSS developers make use of “boundary organisations,” such as consortia or foundations [50, 66]. These institutions provide the institutional infrastructure that limits the influence of a single company [67] and allows different interest groups, such as volunteers and for-profit companies, to collaborate together in OSS projects [24]. Such organisations enable OSS developers to “negotiate their relationship with companies and, when necessary, to defend their commons-based resources from unwanted influence” [40].
While the private sector has historically been the largest funder of OSS, the public sector has played a key role in funding OSS since the origins of the World Wide Web [68]. Governments support OSS through a range of channels, including research grants [69, 70] and dedicated funds, such as the Open Technology Fund in the USA, the EC’s Next Generation Initiative, or the STF in Germany [8]. Lately, OSS has moved up the priority list for governments due to concerns about software security [2] and digital sovereignty [1, 71]. For example, the EC’s OSS strategy argues that OSS can give “Europe a chance to create and maintain its own, independent digital approach and stay in control of its processes” [72] . In October 2022, the German government set up the STF to “invest in [the] digital commons, volunteer communities, and the open source ecosystem,” reasoning that thriving OSS ecosystems are key to digital sovereignty, securing software, economic growth, and digital democracy [7]. The same year, the European Working Team on Digital Commons, comprising 19 EU member states and the EC, underlined the potential of OSS to strengthen digital sovereignty by minimising vendor lock-in, increasing the bargaining power of states, and offering greater design control in the public sector [6].
@@
The public sector’s embrace of OSS to build digital public services is a recognised strategy to stimulate competition in domestic software markets, which might otherwise be monopolised by industry giants, especially Big Technology companies [73]. For example, it has been estimated that the Circulaire 5608, a French law requiring government agencies to favour OSS when procuring software, led to an increase of nearly 600,000 OSS contributions from French developers per year as well as yearly increases in the national competitiveness of the French IT market [3]. Beyond governmental adoption of OSS, government investments in OSS are often targeted at stimulating national competitiveness. For example, in June 2023, President Emmanuel Macron of France announced a €40 million fund to create a “digital commons” to support the development of open source large language models to support domestic startups and to challenge the dominance of Big Technology companies in the AI industry [74].
While the recent surge in governmental interest in funding OSS has broadly been welcomed by practitioners, it remains an open question regarding how OSS developers view its relative benefits and drawbacks compared to other types of funding, as well as the impact thereof on OSS development and maintenance. As governmental interest and involvement in funding OSS development is currently on the rise, it is timely to examine this question and to offer practical insights from the perspective of OSS developers and maintainers. With this context in mind, this paper examines the role of public and private funding in the scikit-learn project, a widely used Python library for ML, and how its maintainers view the relative benefits and drawbacks of their mixed, public-private funding model.
@@ 3.1 Research aims
This paper investigates the emerging role that governments are playing in funding OSS development and maintenance, and how OSS developers view the relative merits and drawbacks of governmental funding through a case study on the scikit-learn project. In software engineering research, a case study “is an empirical enquiry that draws on multiple sources of evidence to investigate one instance (or a small number of instances) of a contemporary software engineering phenomenon within its real-life context” [75]. Single case studies are suitable when one seeks to provide in-depth insights, generate new hypotheses, or develop theories about a critical, extreme, unusual, or revelatory phenomenon [76]. The scikit-learn project was selected as a revelatory case that could provide in-depth insights into the emergent role of the French government in supporting this OSS project and how its maintainers view the relative merits and drawbacks of the government’s funding compared to the other funding sources that have supported the project to date.
To this end, the study focuses on the following research questions:
  • RQ1: What are the interests of the public and private funders of the scikit-learn project, and how have they aligned or conflicted with the interests of the scikit-learn maintainers?
  • RQ2: How do the maintainers of the scikit-learn project view the relative benefits and drawbacks of public and private funding for OSS development?
  • RQ3: What lessons can be drawn from the scikit-learn project’s funding model for OSS practitioners?

### v4

@@ Abstract
Governments are increasingly funding open source software (OSS) development to support software security, digital sovereignty, and national competitiveness in science and innovation, amongst others. However, little is known about how OSS developers evaluate the relative benefits and drawbacks of emergent governmental funding for OSS. This paper explores this question through a case study on scikit-learn, a popular Python library for machine learning, which has been funded by public research grants, commercial sponsorship, micro-donations, and a €32 million grant announced in France’s artificial intelligence strategy. Through 25 interviews with scikit-learn’s maintainers and funders, this study makes two key contributions to research and practice. First, it contributes novel empirical findings on the effective design and implementation of a public-private funding model in an OSS project, as well as how the maintainers of scikit-learn have designed and employed governance protocols to balance the diverse interests of their funders and to safeguard their community ethos. Second, it offers practical lessons on funding in community-led OSS projects and makes recommendations to practitioners. The paper concludes with a discussion of the key recommendations.
Keywords Open source software, OSS funding models, machine learning, scikit-learn, OSS sustainability
@@ 1 Introduction
Open source software (OSS) has moved up the priority list of governments in light of concerns about digital sovereignty [1], software security [2], and national competitiveness in science and innovation [3], amongst others. The disclosure of the log4j vulnerability in November 2021 is widely seen as a turning point that mobilised national governments, from the United States of America (USA) to Germany, to mobilise public funds for the development and maintenance of OSS [4]. Increasing governmental involvement in funding OSS has been generally well-received by OSS developers. However, little is known about how they view the relative benefits and drawbacks of governmental funding compared to other types of funding. This paper explores this question through a case study on scikit-learn, a Python library for machine learning (ML), which is described as “the Swiss army knife of ML” due to its widespread use [5]. Its sustainability as a community-led project in the industry-dominated field of artificial intelligence (AI) [6] is partly attributed to its mixed funding model, including public research grants, commercial sponsorship, community donations, and the institutional backing from the French Institute for Research in Computer Science and Automation (Inria). In November 2021, in its national AI strategy, la “Stratégie IA”, the French government announced a €32 million grant to support the expansion of scikit-learn [7], marking a significant expansion of public funding for scikit-learn.
Through 25 interviews with scikit-learn’s maintainers and funders, this paper investigates the historical role of public and private funding in supporting the community-led OSS project. It explores the diverse interests of scikit-learn’s funders, and how they have aligned or conflicted with those of the maintainers (RQ1). The study also examines how the maintainers view the relative benefits and drawbacks of public and private funding (RQ2). Finally, it discusses practical lessons from scikit-learn’s funding model for OSS practitioners (RQ3). This paper makes two key contributions to research and practice. First, it provides novel empirical findings on the design and implementation of a public-private funding model in an OSS project. Furthermore, it sheds light on how the maintainers employed governance protocols to balance the diverse interests of their funders and to safeguard their community ethos. Second, it offers practical lessons on funding in community-led OSS projects. The paper concludes with a discussion of key recommendations to OSS developer communities, companies, and governments.
@@
The paper has the following structure. First, it discusses prior work on the role of funding in supporting the sustainability of OSS projects. Second, it presents the study design and the case of scikit-learn. Then, it presents the results and discusses their implications for research and practice. The paper concludes with a discussion of the key recommendations to practitioners and suggestions for future research directions on OSS funding models.
@@ 2.1 The Role of Funding in OSS Sustainability
The adoption of OSS is ubiquitous, estimated to be used in 96% of code bases [8] and constituting up to 90% of software stacks [9]. As a result, OSS is considered a critical component of the digital infrastructure that underpins modern society and the digital economy [10, 11, 12]. Given its importance, the health and sustainability of OSS projects is a concern for a diverse range of stakeholders. However, the health and sustainability of OSS projects is a complex and multi-faceted issue [13, 14, 15]. A range of factors are relevant, including but not limited to open governance and strong leadership [16, 17, 18, 19]; maintainer best practices, such as timely responses [20], mentoring contributors [21, 22], and recommending tasks [23]; a steady inflow of new contributors [24, 25, 22]; and diversity within the community [26]. Amongst these factors, the financial health of a project plays a crucial, but hitherto understudied, role in the sustainability of OSS projects, determining the extent to which maintenance is funded or reliant on volunteers [27].
@@
However, money has long been a contentious subject amongst OSS developers. On the one hand, many developers take pride in the origins of the OSS movement as a social movement [28], describing it as a “programmers’ paradise” for “geeks”, “hackers”, and “hobbyists”, who do not contribute to OSS for money [29, 30]. Developers cite their political ideals [31], altruism [32], or their passion for spotting and quashing bugs in software [33] as intrinsic incentives for contributing to OSS. Furthermore, many developers complain that “money ruins everything” [34]. On the other hand, a plethora of OSS projects are unfunded and maintained by overstretched volunteers, who often struggle to keep up with the workload required to keep projects running and communities thriving [27, 35, 20].
@@
The consequences of systemic issues of free-riding and under-investment in OSS are highlighted when developers spot major vulnerabilities that pose significant threats to digital systems in governments, critical infrastructure, and companies’ products and services. In particular, the discovery of vulnerabilities, such as the heartbleed bug in April 2014 or log4j vulnerability in December 2021, which was described as the “worst security problem in a generation” [36], shocked industry and governments alike. It highlighted the need for cross-sector stakeholders, including the public and private sectors, to fund the developer communities who build and maintain our open source digital infrastructure [4]. As Benjamin Bikinbine has argued, there is an urgent need “not just [for] investment in institutions, organisations, technologies, or innovations, but long-term and sustainable investment in the true source of their value: people” [37].
2.2 Private Funding for OSS Development
The private sector has been the largest funder of OSS to date, funding OSS developers and projects in both direct and indirect ways. The impact of commercial investments in OSS are gauged to be significant. According to one estimate, companies in the European Union (EU) invested around €1 billion in OSS in 2018, which generated between €65-95 billion for the EU’s GDP that year [38]. Direct funding methods include sponsorship of OSS foundations and consortia, which operate on a fee-paying membership model [39]. Foundations and consortia are known to play a critical role as “boundary organisations”, which through their open governance protocols facilitate collaboration between diverse volunteers and companies [16]. However, according to a survey of 369 OSS developers, only 12% of respondents reported that organisational funding for OSS via a foundation, a consortium, or an independent legal entity was either extremely or very useful to their development activity, while 5% reported it was ineffective [40].
@@ 2.2 Private Funding for OSS Development
Recently, companies have begun to establish FOSS Contributor Funds, which allow employees to nominate and vote on OSS projects that should receive funding [41], and represent a more “democratic” way for companies to fund the OSS projects that they use [42]. A shortcoming is that these funds typically award up to \$10,000 over a one-year time frame. Alyssa Wright from Bloomberg’s FOSS Contributor Fund [43] has acknowledged this shortcoming, arguing that “true sustainability comes from a multi-sided commitment” and “while financial support is important, real engagement through contribution and dialogue is as crucial to the ecosystem’s survival as writing a check”.
@@
Other ways companies support OSS projects by sponsoring OSS developers or by allowing their employees to contribute to OSS projects at work [44, 45]. Sponsoring developers is a recognised strategy that companies employ to influence projects [46, 47] and to improve their reputation as OSS patrons, which in turn is considered to aid the recruitment of developers [48, 49, 50]. In the mid-2000s, it was estimated that 40% of OSS contributors were paid [51]. We should expect this number to be much higher today, with companies like Google, Microsoft, and Amazon accounting for a rapidly growing number of OSS developers and contributions to OSS projects [52]. In addition, companies invest in projects through donations and by joining project steering committees [53].
While OSS projects benefit from commercial participation, maintainers seek to manage the influence or the dominance of individual companies [54]. A popular strategy is to establish or transfer projects to foundations [39, 55], which through open governance protocols limit the dominance of single companies [56] and facilitate collaboration between diverse contributors [16, 57]. Crucially, these organisations enable OSS developers to “negotiate their relationship with companies and, when necessary, to defend their commons-based resources from unwanted influence” [37].
2.3 Micro-Donations for OSS Development
Individuals or organisations support OSS developers and project by making micro-donations to individual developers and OSS projects. For instance, the GitHub Sponsors scheme, launched in 2019, allows users to make one-off or recurring payments to individuals or organisations with sponsored profiles [58]. According to prior work, common incentives for donations via GitHub Sponsors include expressing appreciation, recognising a developer’s work, and encouraging future contributions [30]. By comparison, common incentives for maintainers to accept donations are to receive recognition for their work or to provide a side-income; while those who do not participate in GitHub Sponsors state that they do not need to be sponsored or that they do not contribute to OSS for money [30]. Furthermore, tther work shows that the amount that individuals donate is associated with the length of their use of an OSS [59]. The impact of micro-donations is still contested. On the one hand, they have been shown to shorten response times to issues [60, 61] and to increase maintenance-related activities [62]. On the other hand, overall the impact appears trivial [63] and donations tend to be ad-hoc, small, or legally complicated for individuals to receive [64].
2.4 Public Funding for OSS Development
The public sector has indirectly or directly funded OSS since the origins of the World Wide Web [65] through research grants for (scientific) OSS [66, 67, 68] as well as bespoke OSS funding bodies, such as the Open Technology Fund in the USA, the Next Generation Initiative at the European Commission, or the Sovereign Tech Fund (STF) in Germany [4]. Furthermore, OSS has recently moved up the priority list of governments in light of concerns about the security of OSS [69, 2]; its role as critical digital infrastructure [10, 11, 12]; and its importance for digital sovereignty [1, 70] and national competitiveness in science and innovation [3]. In Europe, the concern for digital sovereignty is particularly salient. For example, the European Commission’s OSS strategy argues that OSS can give “Europe a chance to create and maintain its own, independent digital approach and stay in control of its processes” [71].
@@ 2.4 Public Funding for OSS Development
Public sector support for OSS is also a recognised strategy to stimulate competition in domestic software markets, which might otherwise be monopolised by industry giants [72]. For example, it has been estimated that the Circulaire 5608, a French law requiring government agencies to favour OSS when procuring software, led to an increase of nearly 600,000 OSS contributions from French developers per year as well as yearly increases in the national competitiveness of the French IT market [3]. Beyond governmental adoption of OSS, government investments in OSS are often targeted at stimulating national competitiveness. For example, in June 2023, President Emmanuel Macron of France announced a €40 million fund to create a “digital commons” to support the development of open source large language models to support domestic startups and to challenge the dominance of Big Technology companies in the AI industry [73].
The discovery of the log4j vulnerability in November 2021 marked a turning point for governmental interest and involvement in funding OSS development. In light of the magnitude of the security concerns [69], governments rapidly mobilised to address the threats posed by OSS vulnerabilities to their critical infrastructure and digital economies. For example, in January 2022, the White House convened cross-sector stakeholders, including government agencies, Big Technology companies, and OSS foundations, to discuss the urgent need to improve the security of OSS, recognising its indispensable role as digital infrastructure whose maintenance largely depends on volunteers [74]. Across the pond, in June 2022, the newly formed European Working Team on the Digital Commons, comprising representatives from 19 member states of the EU and the European Commission, published its inaugural report, highlighting the importance of funding OSS for bolstering the digital sovereignty of Europe [75]. In October 2022, Germany established the STF, allocating a budget of €11.5 million for its first year alone to fund OSS maintenance in the interest of digital sovereignty, software security, innovation, and digital democracy [76]. These developments underscore the growing recognition amongst policymakers of the public value of funding OSS development and maintenance [4].
While there is a growing recognition of the need for sustainable funding for OSS by governments, little is known about how OSS developers evaluate the relative merits and drawbacks of governmental involvement in funding OSS development compared to other types of funding that have primarily supported OSS development so far, or the impact thereof on OSS development and maintenance. This study addresses this gap by examining the case of scikit-learn and how its maintainers view the benefits and drawbacks of its public and private funders. By focusing on the perspectives of OSS developers, this study contributes practical insights on this timely question in the debate about OSS sustainability.
@@ 3.1 Research aims
This paper investigates the emerging role that governments are playing in funding OSS development and maintenance, and how OSS developers view the relative merits and drawbacks of governmental funding through a case study on the scikit-learn project. In software engineering research, a case study “is an empirical enquiry that draws on multiple sources of evidence to investigate one instance (or a small number of instances) of a contemporary software engineering phenomenon within its real-life context” [77]. Single case studies are suitable when one seeks to provide in-depth insights, generate new hypotheses, or develop theories about a critical, extreme, unusual, or revelatory phenomenon [78]. The scikit-learn project was selected as an exploratory case that could provide in-depth insights into the role of the French government in supporting this OSS project and how its maintainers view the relative merits and drawbacks of the government’s funding compared to the other funding sources that have supported the project to date.
To this end, the study focuses on the following research questions (RQ):
  • RQ1: What are the interests of the public and private funders of scikit-learn, and how have they aligned or conflicted with the interests of the scikit-learn maintainers?
  • RQ2: How do the maintainers of scikit-learn view the relative benefits and drawbacks of public and private funding for their OSS project?
  • RQ3: What practical lessons can be drawn from scikit-learn’s funding model for OSS practitioners?